# Model-Agnostic Policy Explanations with Large Language Models

### v1

@@ Abstract
Intelligent agents, such as robots, are increasingly deployed in real-world, human-centric environments. To foster appropriate human trust and meet legal and ethical standards, these agents must be able to explain their behavior. However, state-of-the-art agents are typically driven by black-box models like deep neural networks, limiting their interpretability. We propose a method for generating natural language explanations of agent behavior based only on observed states and actions – without access to the agent’s underlying model. Our approach learns a locally interpretable surrogate model of the agent’s behavior from observations, which then guides a large language model to generate plausible explanations with minimal hallucination. Empirical results show that our method produces explanations that are more comprehensible and correct than those from baselines, as judged by both language models and human evaluators. Furthermore, we find that participants in a user study more accurately predicted the agent’s future actions when given our explanations, suggesting improved understanding of agent behavior.
@@ 1 Introduction
Rapid advances in artificial intelligence and machine learning have led to an increase in the deployment of robots and other embodied agents in human-centric and safety-critical settings (Sun et al., 2020; Fatima & Pasha, 2017; Li et al., 2023). As such, it is vital that practitioners – who may be laypeople that lack domain expertise or knowledge of machine learning – are able to query such agents for explanations regarding why a particular prediction has been made – broadly referred to as explainable AI (Amir et al., 2019; Wells & Bednarz, ; Gunning et al., 2019). While progress has been made in this area, prior works tend to focus on explaining agent behavior in terms of rules (Johnson, 1994), vision-based cues (Cruz & Igarashi, 2021; Mishra et al., 2022), semantic concepts (Zabounidis et al., 2023), or trajectories (Guo et al., 2021). However, it has been shown that laypeople benefit from natural language explanations (Mariotti et al., 2020; Alonso et al., 2017) since they do not require specialized knowledge to understand (Wang et al., 2019), leverage human affinity for verbal communication, and increase trust under uncertainty (Gkatzia et al., 2016).
@@ 4 Quantitative Results and Analysis
Llama-2 Path (Ours) 0.46 0.50 0.54 0.56 0.52
@@
Starling Path (Ours) 0.68 0.67 0.68 0.71 0.64
@@
Table 1: Ratio in which the explanation generated by each method is ranked as best according to a given set of criteria over 100 samples. Explanations are generated with multiple language models over different environments. Rankings are automatically generated by an LLM according to the criteria detailed in Sec. 4.1. Values are bolded if they are significantly better than all other methods for a given model according to a paired $t$-test ($p$-value $<0.05$).
@@
Models: In order to evaluate the effect our method has on various large language models, we generate explanations using four popular models: GPT-4 (OpenAI, 2023), GPT-3.5 (Brown et al., 2020), Starling (Zhu et al., 2023), and LLama 2 (Touvron et al., 2023).
@@ 4.1 Evaluating Comprehensibility and Accuracy
USAR Rescue Path (Ours) 0.70 0.75 0.75 1.00 1.00 1.00 0.90 0.85 0.85
@@
USAR Search Path (Ours) 0.90 0.75 0.25 1.00 1.00 0.70 0.90 0.80 0.35
@@
We first evaluate how our approach affects the quality of generated explanations. Unlike works that evaluate natural language explanations in domains such as natural language inference, to the best of our knowledge, there are no datasets consisting of high-quality explanations generated over agent behavior. Subsequently, we create annotated datasets by sampling agent behaviors from heuristic policies in each environment. Using a heuristic policy allows us to access the internal beliefs and intents of the agent and programmatically generate ground truth explanations. We generate explanations for each observation in this dataset using the methods described previously and automatically rank them using an LLM (GPT-4o) (Zhu et al., 2023) according to five criteria: informativeness which measures whether an explanation contains sufficient information to understand an action; correctness which measures whether an explanation contains accurate information; and whether an explanation identifies an agent’s strategy, goal, and goal category. Definitions for each criteria can be found in Appendix A.2. Our results are shown in Table 1 where we observe the following.
@@ 5.1 Evaluating Helpfulness through Action Prediction Performance (H1)
Using binary logistic regression with controls, we find that H1.1 ($coef=+2.650$, $SE=0.436$, $oddsratio=14.15$, $p<0.001$), H1.2 ($coef=+2.898$, $SE=0.617$, $oddsratio=18.14$, $p<0.001$), and H1.3 ($coef=+1.390$, $SE=0.508$, $oddsratio=4.014$, $p<0.01$) are all supported with significance (see Fig. 4). That is, participants are significantly more accurate in predicting the agent’s next action when provided with explanations generated by our method, compared to those from State, Template, or no explanation at all. A McNemar’s test shows no significant difference in accuracy between Path and Human explanations, supporting the hypothesis H1.4 $(p>0.05)$ that our explanations perform on par with those written by human experts.
@@ 5.2 Evaluating Helpfulness through Participant Preference (H2)
Using Wilcoxon signed-rank tests, we find that H2.2 ($W=3016.00$, $n=81$, $r=+0.709$, $p<0.001$) is supported with significance: participants show a clear preference for explanations generated by our method over those from Template. In contrast, H2.1 ($W=1778.00$, $n=81$, $r=+0.061$, $p>0.05$) is not supported, indicating no significant difference in preference between our method and State. H2.3 ($W=741.00$, $n=74$, $r=-0.405$, $p<0.001$) is also not supported, with participants significantly preferring Human explanations over those generated by our method. See Fig. 5 for full results.
@@ 5.3 Evaluating Hallucination (H3, H4)
Consistent with LLM-based action prediction results in Sec. 4.1, we find that participant action prediction accuracy is not significantly affected by perceived hallucination. Specifically, H3.1 is only supported in Path vs Human comparisons ($p<0.001$) while H3.2 ($p>0.05$) is not supported. Regarding participant preferences, we observe that H4.1 and H4.2 are only supported in the Path-vs.-Template comparison ($p<0.01$ and $p<0.05$), indicating that hallucination does not significantly influence preference in general. Interestingly, subjectively perceived hallucinations in Template strongly decreases preference for Path (p <.01).
To test whether this phenomenon is tied to participant’s inability to detect hallucinations, we further perform another binary logistic regression with controls using hallucination labels from both tasks. Results suggest that there is no significance ($coef=+0.482,SE=0.942,oddsratio=1.620,p>0.05$) between objectively present and subjectively perceived hallucination, indicating participants were unable to detect hallucinated information.
@@ 7 Limitations
Our work currently has several limiting assumptions. First, we assume that the task for which explanations are generated falls within the pre-trained LLM’s training data distribution. That is, we don’t expect this approach to work if it requires specialized knowledge, e.g. neurosurgical expertise, if that is lacking from the LLM’s training data. Along these lines, we conjecture that our approach is better-suited for high-level explanations, e.g. the robot went to the kitchen, rather than low-level granular explanations, e.g. the robot applied a trajectory of velocity commands.
Second, we currently leverage a decision tree as a behavior representation, which requires structured inputs and precludes operating over dense, unstructured inputs such as images. However, we believe this limitation can be overcome by employing alternative models for behavior representations, e.g. feature extractors or differentiable decision trees.
Lastly, as with all post-hoc explainability methods, there is no guarantee that the generated explanations represent the actual decision making process of the underlying agent model. However, in practice an explanation only needs to be faithful enough to be useful, which in turn depends on the purpose and audience. For example, humans ascribe mental models to explain the behavior of other humans, pets, or even robots so that we may plan our own actions. These models may not be fully accurate to the underlying reasoning process, but if they roughly capture high-level behavior then they are still useful. We view post-hoc explanations such as the ones presented in this work in the same light.
Separately, our participant study in Sec. 5 has two notable limitations. First, our sample population of university students may not adequately represent a larger population for typical people. Second, for participant helpfulness rating, we used a single Likert item rather than a full Likert scale. While a full scale would generally yield more reliable assessments, we chose a single item to reduce the risk of survey fatigue.
@@ A.1.1 Prompt Sensitivity Analysis
Table 3: Ratio in which the explanation generated by each method is ranked as best according to a given set of criteria over 100 samples. Explanations are generated with altered prompts adhereing to task id 280 and 904 Sclar et al. (2023). Rankings are automatically generated by an LLM according to the criteria detailed in Sec. 4.1. Values are bolded if they are significantly better than all other methods for a given model according to a paired $t$-test ($p$-value $<0.05$).
@@ A.1.2 Explanation Quality for USAR Fixed
Fixed Path (Ours) 0.80 0.40 0.25
@@ A.5 Participant Study
We conducted a participant study approved by the Georgia Institute of Technology Institutional Review Board (Protocol No. IRB2025-83). This study evaluates three behavior explanation methods (Path, State, and Template) based on helpfulness, prediction accuracy, and hallucination detection. Each participant will complete two task types: comparison (Type A) and action prediction (Type B), using 8 cases split evenly. In Type A tasks, participants compare explanations pairwise, rank their helpfulness, and note if they seem hallucinated. In Type B tasks, participants predict the AI agent’s next move, first without and then with explanations from each method, and note if they seem hallucinated. Maps and agent icons will be altered to avoid bias. Participants will provide demographic data, receive an introduction to the grid world environment, and complete tasks in the form of multiple choice questions. For each multiple-choice question, visualizations in GIF format that show the agent’s trajectory will be provided. Results will be analyzed using statistical analysis to determine the most effective explanation method.
@@ A.5.1 Statistical Results
Here we describe our statistical approaches, and report all relevant statistical values for our hypothesis in Sec. 5.
@@
H1: Participants predicted the next action significantly more accurately with explanations generated by Path compared to State (H1.1), Template (H1.2), and no explanation (H1.3), with no significant difference in accuracy compared to Human explanations (H1.4).
  • H1.1 (Path >No Explanation): Logistic regression results revealed a significant positive effect of the Path explanation compared to no explanation on prediction accuracy ($coef=+2.650,SE=0.436,oddsratio=14.15,p<0.001$), supporting H1.1.
  • H1.2 (Path >State): Path explanations led to significantly higher prediction accuracy than State explanations ($coef=+2.898,SE=0.617,oddsratio=18.14,p<0.001$), supporting H1.2.
  • H1.3 (Path >Template): Participants were also significantly more accurate with Path explanations compared to Template explanations ($coef=+1.390,SE=0.508,oddsratio=4.014,p=0.006$), supporting H1.3.
  • H1.4 (Path = Human): With 16 discordant pairs, we applied the exact binomial version of the test, which revealed no significant difference ($p=0.455$) between the two explanation types. Thus, H1.4 is supported.
H3: Participants predicted the next action less accurately when explanations contained hallucinations, whether these were subjectively perceived (H3.1) or objectively present (H3.2).
[…]
  • H3.2 (Objective Hallucinations): Objectively present hallucinations did not significantly affect prediction accuracy in Path vs. No explanation ($coef=+0.031,SE=0.391,oddsratio=1.03,p=0.936$), Path vs. State ($coef=-0.498,SE=0.522,oddsratio=0.61,p=0.340$), Path vs. Template ($coef=+0.451,SE=0.648,oddsratio=1.57,p=0.486$), and Path vs. Human ($coef=+0.234,SE=0.664,oddsratio=1.26,p=0.724$); thus, H3.2 is not supported.
@@
H2: Participants prefer explanations generated by Path compared to State (H2.1) and Template (H2.2), with no significant difference in preference compared to Human explanations (H2.3).
  • H2.1 (Path >State): A one-sided Wilcoxon signed-rank test shows that participants do not significantly prefer Path over State ($W=1778.00,n=81,r=+0.061,p=0.2839$). H2.1 is not supported.
  • H2.2 (Path >Template): A one-sided Wilcoxon signed-rank test shows that participants significantly prefer Path over Template ($W=3016.00,n=81,r=+0.709,p<0.001$). H2.2 is supported.
  • H2.3 (Path = Human): A two-sided Wilcoxon signed-rank test revealed a significant preference for Human explanations over Path explanations ($W=741.00,n=74,r=-0.405,p<0.001$). H2.3 is not supported.
H4: Explanations containing hallucinations—both subjectively perceived (H4.1) and objectively present (H4.2)—receive lower preference ratings than those without hallucinations.
@@
    • Path vs. Template: perceived hallucinations in both explanations significantly decreased preference for Path ($coef=-1.763,SE=0.600,oddsratio=0.172,p=0.003$ for Path; $coef=-1.758,SE=0.539,oddsratio=0.172,p=0.001$ for Template).
    • Path vs. State: non-significant positive effects ($coef=+0.390,SE=0.459,oddsratio=1.477,p=0.396$ for Path; $coef=+0.657,SE=0.465,oddsratio=1.928,p=0.1575$ for State).
    • Path vs. Human: mixed, non-significant effects ($coef=+0.662,SE=0.440,oddsratio=1.939,p=0.133$ for Path; $coef=-0.350,SE=0.432,oddsratio=0.705,p=0.417$ for State).
@@
    • Path vs. Template: Ground truth hallucinations in Path significantly decreases participant preference for Path ($coef=-1.049,SE=0.503,oddsratio=0.350,p=0.037$). Template doesn’t contain hallucinations; not applicable for study.
    • Path vs. State: Ground truth hallucinations in both explanations do not significantly affect the preferences of the participants ($coef=+0.515,SE=0.468,oddsratio=1.674,p=0.272$ for Path; $coef=+0.387,SE=0.478,oddsratio=1.473,p=0.418$ for State).
    • Path vs. Human: Ground truth hallucinations in Path do not significantly affect the preferences of the participants ($coef=+0.096,SE=0.429,oddsratio=1.101,p=0.823$ for Path). Human doesn’t contain hallucinations; not applicable for study.
@@
Additional Statistics We perform a binary logistic regression with controls to test whether subjectively perceived hallucinations are significantly correlated with objectively present hallucination, and find that the hypothesis is rejected $coef=+0.482,SE=0.942,oddsratio=1.620,p=0.609$.

### v2

@@ Abstract
Intelligent agents, such as robots, are increasingly deployed in real-world, human-centric environments. To foster appropriate human trust and meet legal and ethical standards, these agents must be able to explain their behavior. However, state-of-the-art agents are typically driven by black-box models like deep neural networks, limiting their interpretability. We propose a method for generating natural language explanations of agent behavior based only on observed states and actions – without access to the agent’s underlying model. Our approach learns a locally interpretable surrogate model of the agent’s behavior from observations, which then guides a large language model to generate plausible explanations with minimal hallucination. Empirical results show that our method produces explanations that are more comprehensible and correct than those from baselines, as judged by both language models and human evaluators. Furthermore, we find that participants in a user study more accurately predicted the agent’s future actions when given our explanations, suggesting improved understanding of agent behavior. Importantly, we show that participants are unable to detect hallucinations in explanations, underscoring the need for explainability methods that minimize hallucinations by design.
@@ 1 Introduction
Rapid advances in artificial intelligence and machine learning have led to an increase in the deployment of robots and other embodied agents in human-centric and safety-critical settings (Sun et al., 2020; Fatima & Pasha, 2017; Li et al., 2023). As such, it is vital that practitioners – who may be laypeople that lack domain expertise or knowledge of machine learning – are able to query such agents for explanations regarding why a particular prediction has been made – broadly referred to as explainable AI (Amir et al., 2019; Wells & Bednarz, 2021; Gunning et al., 2019). While progress has been made in this area, prior works tend to focus on explaining agent behavior in terms of rules (Johnson, 1994), vision-based cues (Cruz & Igarashi, 2021; Mishra et al., 2022), semantic concepts (Zabounidis et al., 2023), or trajectories (Guo et al., 2021). However, it has been shown that laypeople benefit from natural language explanations (Mariotti et al., 2020; Alonso et al., 2017) since they do not require specialized knowledge to understand (Wang et al., 2019), leverage human affinity for verbal communication, and increase trust under uncertainty (Gkatzia et al., 2016).
@@ 4 Quantitative Results and Analysis
Llama-2
Path (Ours) 0.46 0.50 0.54 0.56 0.52
@@
Starling
Path (Ours) 0.68 0.67 0.68 0.71 0.64
@@
Table 1: Ratio in which the explanation generated by each method is ranked as best according to a given set of criteria over 100 samples. Explanations are generated with multiple language models over different environments. Rankings are automatically generated by an LLM according to the criteria detailed in Sec. 4.1. Values are bolded if they are significantly better than all other methods for a given model according to a paired $t$-test ($p$-value <0.05<0.05).
@@
Models: In order to evaluate the effect our method has on various large language models, we generate explanations using four popular models: GPT-4 (OpenAI, 2023), GPT-3.5 (Brown et al., 2020), Starling (Zhu et al., 2024), and LLama 2 (Touvron et al., 2023).
@@ 4.1 Evaluating Comprehensibility and Accuracy
USAR
Rescue
Path (Ours) 0.70 0.75 0.75 1.00 1.00 1.00 0.90 0.85 0.85
@@
USAR
Search
Path (Ours) 0.90 0.75 0.25 1.00 1.00 0.70 0.90 0.80 0.35
@@
We first evaluate how our approach affects the quality of generated explanations. Unlike works that evaluate natural language explanations in domains such as natural language inference, to the best of our knowledge, there are no datasets consisting of high-quality explanations generated over agent behavior. Subsequently, we create annotated datasets by sampling agent behaviors from heuristic policies in each environment. Using a heuristic policy allows us to access the internal beliefs and intents of the agent and programmatically generate ground truth explanations. We generate explanations for each observation in this dataset using the methods described previously and automatically rank them using an LLM (GPT-4o) (Zhu et al., 2024) according to five criteria: informativeness which measures whether an explanation contains sufficient information to understand an action; correctness which measures whether an explanation contains accurate information; and whether an explanation identifies an agent’s strategy, goal, and goal category. Definitions for each criteria can be found in Appendix A.2. Our results are shown in Table 1 where we observe the following.
@@ 5.1 Evaluating Helpfulness through Action Prediction Performance (H1)
Using binary logistic regression with controls, we find that H1.1 ($coef=+2.650$, $p<0.001$), H1.2 ($coef=+2.898$, $p<0.001$), and H1.3 ($coef=+1.390$, $p<0.01$) are all supported with significance (see Fig. 4). That is, participants are significantly more accurate in predicting the agent’s next action when provided with explanations generated by our method, compared to those from State, Template, or no explanation at all. A McNemar’s test shows no significant difference in accuracy between Path and Human explanations, supporting the hypothesis H1.4 $(p>0.05)$ that our explanations perform on par with those written by human experts.
@@ 5.2 Evaluating Helpfulness through Participant Preference (H2)
Using Wilcoxon signed-rank tests, we find that H2.2 ($W=3016.00$, $p<0.001$) is supported with significance: participants show a clear preference for explanations generated by our method over those from Template. In contrast, H2.1 ($W=1778.00$, $p>0.05$) is not supported, indicating no significant difference in preference between our method and State. H2.3 ($W=741.00$, $p<0.001$) is also not supported, with participants significantly preferring Human explanations over those generated by our method. See Fig. 5 for full results.
@@ 5.3 Evaluating Hallucination (H3, H4)
Consistent with LLM-based action prediction results in Sec. 4.1, we find that participant action prediction accuracy is not significantly affected by perceived hallucination. Specifically, H3.1 is only supported in Path vs Human comparisons (coef=1.713,p<0.001coef=-1.713,p<0.001) while H3.2 ($p>0.05$) is not supported. Regarding participant preferences, we observe that H4.1 and H4.2 are only supported in the Path-vs.-Template comparison (coef=1.763,p=0.01coef=-1.763,p=0.01 and coef=1.049,p<0.05coef=-1.049,p<0.05), indicating that hallucination does not significantly influence preference in general. Interestingly, subjectively perceived hallucinations in Template strongly decreases preference for Path (coef=1.758,p<0.01coef=-1.758,p<0.01).
To test whether this phenomenon is tied to participant’s inability to detect hallucinations, we further perform another binary logistic regression with controls using hallucination labels from both tasks. Results suggest that there is no significance ($coef=+0.482,p>0.05$) between objectively present and subjectively perceived hallucination, indicating participants were unable to detect hallucinated information.
@@ 7 Limitations
Our work currently has several limiting assumptions. First, we assume that the task for which explanations are generated falls within the pre-trained LLM’s training data distribution. That is, we don’t expect this approach to work if it requires specialized knowledge, e.g. neurosurgical expertise, if that is lacking from the LLM’s training data. Along these lines, we conjecture that our approach is better-suited for high-level explanations, e.g. the robot went to the kitchen, rather than low-level granular explanations, e.g. the robot applied a trajectory of velocity commands. Second, we currently leverage a decision tree as a behavior representation, which requires structured inputs and precludes operating over dense, unstructured inputs such as images. However, we believe this limitation can be overcome by employing alternative models for behavior representations, e.g. feature extractors or differentiable decision trees. Lastly, as with all post-hoc explainability methods, there is no guarantee that the generated explanations represent the actual decision making process of the underlying agent model. However, in practice an explanation only needs to be faithful enough to be useful, which in turn depends on the purpose and audience. For example, humans ascribe mental models to explain the behavior of other humans, pets, or even robots so that we may plan our own actions. These models may not be fully accurate to the underlying reasoning process, but if they roughly capture high-level behavior then they are still useful. We view post-hoc explanations such as the ones presented in this work in the same light.
In addition, our participant study has two notable limitations. First, our sample population of university students may not be representative of potential users. Second, for participant helpfulness ratings, we used a single Likert item for comparisons—this isn’t ideal and a Likert scale would generally be better. However, we opted for a Likert item because we felt that administering a full Likert scale would result in survey fatigue.
@@ A.1.1 Prompt Sensitivity Analysis
Table 3: Ratio in which the explanation generated by each method is ranked as best according to a given set of criteria over 100 samples. Explanations are generated with altered prompts adhereing to task id 280 and 904 Sclar et al. (2023). Rankings are automatically generated by an LLM according to the criteria detailed in Sec. 4.1. Values are bolded if they are significantly better than all other methods for a given model according to a paired $t$-test ($p$-value <0.05<0.05).
@@ A.1.2 Explanation Quality for USAR Fixed
Fixed
Path (Ours) 0.80 0.40 0.25
@@
A.1.3 Human Preference Alignment
To further understand how well human preferences align with explanation correctness and utility, we conducted additional statistical analyses on our participant data. In examining how preferences align with correctness, we found that participants tend to be overly generous in accepting explanations that did not reflect the current agent’s strategy (as discussed in Section 5.2), and instead appear to rely on other factors when making their decisions. To investigate what these factors might be, we ran a logistic regression to assess whether explanation length influenced preference ratings. We found that longer explanations were significantly less likely to be preferred ($p<0.001$ without controls; marginally significant at $p\approx 0.053$ with controls). This suggests that the relationship between preference and correctness may be confounded or even dominated by differences in explanation length. In practice, we observed that participants tended to select shorter responses, regardless of their fidelity.
As for explanation utility, we note that presenting both explanations side by side made it difficult to isolate the attention participants gave to each. However,logistic regression on total time spent per question found no evidence that longer engagement led to better preference for non-hallucinated explanations (p = 0.43 without controls; p = 0.38 with controls), nor to more accurate hallucination detection (p = 0.74 without controls; p = 1.00 with controls). This supports our finding that participants were not able to reliably detect hallucinations (Sec. 5.3), and further motivates the need for developing methods that inherently produce correct explanations.
@@ A.5 Participant Study
This study evaluates three behavior explanation methods (Path, State, and Template) based on helpfulness, prediction accuracy, and hallucination detection. Each participant will complete two task types: comparison (Type A) and action prediction (Type B), using 8 cases split evenly. In Type A tasks, participants compare explanations pairwise, rank their helpfulness, and note if they seem hallucinated. In Type B tasks, participants predict the AI agent’s next move, first without and then with explanations from each method, and note if they seem hallucinated. Maps and agent icons will be altered to avoid bias. Participants will provide demographic data, receive an introduction to the grid world environment, and complete tasks in the form of multiple choice questions. For each multiple-choice question, visualizations in GIF format that show the agent’s trajectory will be provided. Results will be analyzed using statistical analysis to determine the most effective explanation method.
@@ A.5.1 Statistical Results
Here we describe our statistical approaches and report all relevant statistical values for our hypothesis in Sec. 5 to ensure transparency and reproducibility.
@@
  • H1.1 (Path >No Explanation):(coef=+2.650,SE=0.436,oddsratio=14.15,p<0.001coef=+2.650,SE=0.436,oddsratio=14.15,p<0.001).
  • H1.2 (Path >State): (coef=+2.898,SE=0.617,oddsratio=18.14,p<0.001coef=+2.898,SE=0.617,oddsratio=18.14,p<0.001).
  • H1.3 (Path >Template): (coef=+1.390,SE=0.508,oddsratio=4.014,p=0.006coef=+1.390,SE=0.508,oddsratio=4.014,p=0.006).
  • H1.4 (Path = Human): ($p=0.455$).
  • H3.1 (Subjective Hallucinations):
    • Path vs. Human: (coef=1.713,SE=0.447,OR=0.18,p<.001coef=-1.713,SE=0.447,OR=0.18,p<.001)
    • Path vs. No Explanation: (coef=0.395,SE=0.340,OR=0.67,p=0.245coef=-0.395,SE=0.340,OR=0.67,p=0.245)
    • Path vs. State: (coef=0.918,SE=0.519,OR=0.40,p=0.077coef=-0.918,SE=0.519,OR=0.40,p=0.077)
    • Path vs. Template: (coef=0.692,SE=0.478,OR=0.50,p=0.148coef=-0.692,SE=0.478,OR=0.50,p=0.148)
  • H3.2 (Objective Hallucinations):
    • Path vs. No Explanation: ($coef=+0.031,SE=0.391,OR=1.03,p=0.936$)
    • Path vs. State: (coef=0.498,SE=0.522,OR=0.61,p=0.340coef=-0.498,SE=0.522,OR=0.61,p=0.340)
    • Path vs. Template: ($coef=+0.451,SE=0.648,OR=1.57,p=0.486$)
    • Path vs. Human: ($coef=+0.234,SE=0.664,OR=1.26,p=0.724$)
@@
  • H2.1 (Path >State): ($W=1778.00,n=81,r=+0.061,p=0.284$)
  • H2.2 (Path >Template): ($W=3016.00,n=81,r=+0.709,p<0.001$)
  • H2.3 (Path = Human): (W=741.00,n=74,r=0.405,p<0.001W=741.00,n=74,r=-0.405,p<0.001)
@@
    • Path vs. Template: Path (coef=1.763,SE=0.600,OR=0.172,p=0.003coef=-1.763,SE=0.600,OR=0.172,p=0.003), Template (coef=1.758,SE=0.539,OR=0.172,p=0.001coef=-1.758,SE=0.539,OR=0.172,p=0.001)
    • Path vs. State: Path ($coef=+0.390,SE=0.459,OR=1.48,p=0.396$), State ($coef=+0.657,SE=0.465,OR=1.93,p=0.158$)
    • Path vs. Human: Path ($coef=+0.662,SE=0.440,OR=1.94,p=0.133$), Human (coef=0.350,SE=0.432,OR=0.71,p=0.417coef=-0.350,SE=0.432,OR=0.71,p=0.417)
@@
    • Path vs. Template: Path (coef=1.049,SE=0.503,OR=0.35,p=0.037coef=-1.049,SE=0.503,OR=0.35,p=0.037); Template has no hallucinations.
    • Path vs. State: Path ($coef=+0.515,SE=0.468,OR=1.67,p=0.272$), State ($coef=+0.387,SE=0.478,OR=1.47,p=0.418$)
    • Path vs. Human: Path ($coef=+0.096,SE=0.429,OR=1.10,p=0.823$); Human has no hallucinations.
@@
Additional Statistics We perform a binary logistic regression with controls to test whether subjectively perceived hallucinations are significantly correlated with objectively present hallucination, and find that the hypothesis is rejected coef=+0.482,SE=0.942,oddsratio=1.620,p=0.609coef=+0.482,SE=0.942,oddsratio=1.620,p=0.609.